### v1

# Self-Supervised Multi-Part Articulated Objects Modeling via Deformable Gaussian Splatting and Progressive Primitive Segmentation

Abstract
Articulated objects are ubiquitous in everyday life, and accurate 3D representations of their geometry and motion are critical for numerous applications. However, in the absence of human annotation, existing approaches still struggle to build a unified representation for objects that contain multiple movable parts. We introduce DeGSS, a unified framework that encodes articulated objects as deformable 3D Gaussian fields, embedding geometry, appearance, and motion in one compact representation. Each interaction state is modeled as a smooth deformation of a shared field, and the resulting deformation trajectories guide a progressive coarse-to-fine part segmentation that identifies distinct rigid components, all in an unsupervised manner. The refined field provides a spatially continuous, fully decoupled description of every part, supporting part-level reconstruction and precise modeling of their kinematic relationships. To evaluate generalization and realism, we enlarge the synthetic PartNet-Mobility benchmark and release RS-Art, a real-to-sim dataset that pairs RGB captures with accurately reverse-engineered 3D models. Extensive experiments demonstrate that our method outperforms existing methods in both accuracy and stability.
1 Introduction
Articulated objects, which range from hinged doors and sliding drawers to foldable laptops and many other everyday items with movable parts, are ubiquitous in both physical and virtual environments. Accurate modeling of their geometry and kinematics is indispensable for a wide spectrum of downstream tasks because it empowers robots to manipulate objects reliably [1, 2, 3, 4], enables plausible interactions in AR/VR scenes [5, 6, 7], and provides structured assets for digital twins [8, 9, 10]. These requirements motivate research on unified frameworks that simultaneously recover the geometry, appearance, and motion dynamics of articulated objects at the part level.
Earlier work on articulated object reconstruction can be divided into two strands. The first strand relies on explicit geometric representations and assumes access to clean CAD models with dense part annotations [11, 12], a requirement that limits applicability to real-world sensory data. The second strand adopts continuous implicit shape representations [13] such as signed distance fields to capture the deformation behavior [14, 15, 16, 17]. Their simplified appearance models and coarse geometric priors confine these methods to structurally simple objects with limited diversity.
Recent progress in radiance field rendering, including Neural Radiance Fields (NeRFs) [18] and 3D Gaussian Splatting (3DGS) [19], offers dense photometric self-supervision that jointly captures view-dependent appearance and geometry, providing a prior for articulated object modeling. PARIS [20] first employs a NeRF backbone [18] to analyze two multiview image sets captured in different interaction states and reconstructs separate static and dynamic fields for part-level modeling. Subsequent studies [21, 22] introduce 3DGS [19] to achieve higher rendering fidelity and faster convergence. However, these pipelines assume a single movable part observed in only two interaction states. As a result, each object is represented by one static and one moving component, which confines the framework to simple single-joint objects. Some approaches [23, 24] reconstruct radiance fields for two interaction states, compare the resulting geometry or fit a transform between the states, and then infer a multi-part decomposition. However, they require the number of parts to be known beforehand, and the reliance on exactly two discrete states forfeits any ability to represent continuous motion. In general, the above methods for modeling objects with multiple parts either tackle one joint at a time or assume a predefined part count, and both strategies require explicit correspondences between parts observed in different states. These post hoc correspondence procedures presume strong cross state consistency and prior structural knowledge, conditions that are difficult to satisfy in real scenarios.
Our key insight is that every interaction state of an articulated object can be described as a continuous deformation of a single standard 3D Gaussian field. Specifically, a latent-conditioned deformation network displaces each Gaussian primitive across states, unifying geometry, appearance, and kinematics in one representation and requiring only a handful of multiview image sets for supervision. Differences in deformation trajectories reveal the underlying rigid parts, which we segment through a progressive process: primitives are first clustered by trajectory similarity and then refined with boundary-aware Gaussian splitting that sharpens part interfaces. Drawing on this idea, we introduce DeGSS, a framework that couples Deformable 3D Gaussian Splatting with the Progressive coarse-to-fine Gaussian Primitives Segmentation strategy. Compared to recent methods [20, 23, 21, 22, 24], our approach handles articulated objects with multiple movable parts observed in more than two states, decouples the motion of each part, and smoothly extrapolates to unseen configurations, all without manual annotation.
Moreover, most previous studies [20, 23, 24, 22] test only a single object per category in PartNet-Mobility [1], a synthetic corpus that is too limited for rigorous evaluation and provides little insight into real-world performance. We enlarge the sample size per category in PartNet-Mobility and additionally release RS-Art, a benchmark that pairs multimodal recordings captured for real objects with their corresponding synthetic models, thereby narrowing the gap between simulation and reality.
Our contributions can be summarized as follows:
  • We propose DeGSS, a unified framework that represents articulated objects as deformable Gaussian fields and simultaneously recovers geometry, appearance, and kinematics without additional human priors.
  • We introduce a coarse-to-fine Gaussian segmentation procedure that groups Gaussians by trajectory similarity and refines the clusters, yielding accurate part-aware segmentation.
  • We extend existing benchmarks by enriching PartNet-Mobility and introducing RS-Art, a real-to-sim dataset, and validate the effectiveness and robustness of DeGSS across both benchmarks.
2 Related Work
Articulated object modeling.
Early PointNet pipelines predict part labels and motion from full scans, but depend on CAD-grade geometry and dense supervision, limiting their use in real settings [25, 26, 12]. Implicit SDF methods relax the input requirement, yet still need watertight, pre-aligned frames, a fixed joint budget, and usually only two interaction states [14, 16, 15]. Label-free NeRF approaches such as PARIS fit one static and one dynamic field to a single-joint object, and later 3D Gaussian-splatting variants improve visual fidelity while retaining the same “two-state, one-part” assumption [20, 21, 22]. Recent two-field registration schemes address multi-part objects by reconstructing each state separately and then aligning the fields, but they still require the part count beforehand and assume that the two states are closely aligned in configuration [23, 24]. DeGSS removes these restrictions: a latent-conditioned deformation field models any number of interaction states, while trajectory cues infer the part count automatically, yielding the first fully self-supervised framework for complex multi-part articulation.
Dynamic Gaussian splatting.
Recent work extends static 3D Gaussian splats by making the Gaussian parameters continuous functions of time. Dynamic 3D Gaussians [27] allow each primitive to translate and rotate through a time-conditioned MLP that enforces local rigidity. Four-dimensional Gaussian Splatting [28] embeds Gaussians in a joint space-time volume and renders any time stamp in real time. DynGS [29] introduces a learned trajectory basis that drives per-point motion with a scalar time code, and Explicit-4D Gaussian [30] adopts a similar design while regularizing appearance across frames. Later variants compress or distill these time-dependent fields to improve speed or attach semantics [31]. Although effective for smoothly varying videos, these methods encode dynamics with a single continuous clock, which makes it difficult to model discrete interaction states or multimodal motions where parts jump between configurations. Our method instead learns a conditioned deformation field in which every interaction state is indexed by an individual latent code, enabling explicit state-to-state mapping across arbitrarily many discrete states.
SAM-based 3D segmentation.
Recent efforts lift SAM [32] to neural field segmentation by projecting 2D masks or distilling them into Gaussian features. SA3D [33] propagates a single SAM mask through a trained NeRF to label voxels in three dimensions, and SANeRF HQ [34] improves boundary accuracy by aggregating view masks with density and color cues. Within the Gaussian domain, SAGA [35] learns per Gaussian embeddings under thousands of SAM masks to enable promptable instance grouping, whereas SAGD [36] decomposes boundary Gaussians before fusing multi-view masks for sharper labels. Gaussian Grouping [37] reconstructs scenes with 3DGS and directly lifts SAM masks to obtain object-level segmentation, and SAM3D [38] applies similar ideas to point clouds and meshes. All these methods view segmentation as a static task and rely on externally supplied two-dimensional prompts or per frame masks, so they cannot exploit how primitives move between interaction states. Our framework instead analyses Gaussian deformation trajectories to generate state-aware prompts automatically, then refines them with SAM guided splitting, which yields unmanual-prompt part segmentation.
Figure 1: Overview of the DeGSS pipeline. Solid arrows indicate differentiable modules that participate in the joint optimization, whereas dashed arrows correspond to non-differentiable post-processing stages executed outside the optimization loop.
3 Methodology
This section explains the pipeline shown in Fig. 1. We first review the basics of 3D Gaussian Splatting (Sec. 3.1) and then attach a latent-conditioned deformation network that maps a standard Gaussian field to every interaction state (Sec. 3.2). On the resulting state-specific fields, we detect rigid parts from motion cues (Sec. 3.3) and refine their boundaries with visibility-guided SAM prompting followed by boundary-aware Gaussian splitting (Sec. 3.4). The refined part-aware Gaussian field then supports the multi-task modeling of each articulated object (Sec. 3.5).
3.1 Preliminaries: 3D Gaussian Splatting
3D Gaussian Splatting (3DGS) [19] represents a scene by a finite set of oriented anisotropic Gaussian primitives
$\mathcal{G}=\bigl{\{}g_{i}=(\boldsymbol{\mu}_{i},\;\boldsymbol{\Sigma}_{i},\;\mathbf{c}_{i},\;\alpha_{i})\bigr{\}}_{i=1}^{N},$ (1)
where $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$ denotes the center of the $i$-th Gaussian, $\boldsymbol{\Sigma}_{i}\in\mathbb{R}^{3\times 3}$ its full covariance that encodes both scale and orientation, $\mathbf{c}_{i}\in[0,1]^{3}$ the RGB color, and $\alpha_{i}\in[0,1]$ the opacity.
A perspective camera is modeled by a projection $\Pi:\mathbb{R}^{3}\!\to\!\mathbb{R}^{2}$. Linearization $\Pi$ at $\boldsymbol{\mu}_{i}$ maps the 3D covariance to a 2D screen-space covariance $\mathbf{S}_{i}\in\mathbb{R}^{2\times 2}$ and the mean pixel coordinates $\mathbf{u}_{i}=\Pi(\boldsymbol{\mu}_{i})$. Each primitive therefore induces the density
$\rho_{i}(\mathbf{x})=\alpha_{i}\exp\!\Bigl{(}-\tfrac{1}{2}(\mathbf{x}-\mathbf{u}_{i})^{\!\top}\mathbf{S}_{i}^{-1}(\mathbf{x}-\mathbf{u}_{i})\Bigr{)},\qquad\mathbf{x}\in\mathbb{R}^{2}.$ (2)
With a depth ordering that proceeds from front to back, $\alpha$ compositing yields the pixel radiance
$\mathbf{C}(\mathbf{x})=\sum_{i=1}^{N}w_{i}(\mathbf{x})\,\mathbf{c}_{i},\qquad w_{i}(\mathbf{x})=\rho_{i}(\mathbf{x})\prod_{j<i}\bigl{[}1-\rho_{j}(\mathbf{x})\bigr{]}.$ (3)
All parameters $\{\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i},\mathbf{c}_{i},\alpha_{i}\}_{i=1}^{N}$ are optimized by minimizing a photometric loss between the rendered color $\mathbf{C}(\mathbf{x})$ and the corresponding ground-truth image.
3.2 Deformable Gaussian Splatting
To capture the full continuum of interaction states we train an implicit network, conditioned on a latent code, that warps a standard Gaussian field into state-specific configurations.
Each Gaussian primitive is $g_{i}=(\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i},\mathbf{c}_{i},\alpha_{i})$, with the covariance factorised as
$\boldsymbol{\Sigma}_{i}\;=\;\mathbf{R}\!\left(\mathbf{q}_{i}\right)\,\operatorname{Diag}\!\left(\mathbf{s}_{i}\right)\,\mathbf{R}\!\left(\mathbf{q}_{i}\right)^{\!\top},$ (4)
where $\mathbf{R}(\boldsymbol{q}_{i})$ converts the unit quaternion $\boldsymbol{q}_{i}\in SO(3)$ into its $3\times 3$ rotation matrix, and $\mathrm{Diag}(\boldsymbol{s}_{i})$ places the positive scale vector $\boldsymbol{s}_{i}\in\mathbb{R}_{>0}^{3}$ on the diagonal of a $3\times 3$ matrix. This factorization separates orientation from per-axis stretch, keeping subsequent rotation and scale updates both intuitive and numerically stable.
Given a latent code $\boldsymbol{\alpha}\!\in\!\mathbb{R}^{D}$ that encodes one interaction state, a multi-layer perceptron $f_{\text{def}}$ predicts per-primitive offsets
$(\Delta\boldsymbol{\mu}_{i},\Delta\mathbf{q}_{i},\Delta\mathbf{s}_{i})\;=\;f_{\text{def}}(\boldsymbol{\mu}_{i},\mathbf{q}_{i},\mathbf{s}_{i}\mid\boldsymbol{\alpha}),$ (5)
with $\Delta\boldsymbol{\mu}_{i}\!\in\!\mathbb{R}^{3}$, $\Delta\boldsymbol{s}_{i}\!\in\!\mathbb{R}^{3}$, and an unconstrained four-vector $\Delta\boldsymbol{q}_{i}$ that is normalized to $SO(3)$ by $\Delta\boldsymbol{q}_{i}\leftarrow\Delta\boldsymbol{q}_{i}/\lVert\Delta\boldsymbol{q}_{i}\rVert_{2}$.
Applying the learned offsets converts the standard Gaussian field into the configuration of a particular interaction state $k$:
$\boldsymbol{\mu}_{i}^{(k)}=\boldsymbol{\mu}_{i}+\Delta\boldsymbol{\mu}_{i},\qquad\mathbf{s}_{i}^{(k)}=\mathbf{s}_{i}+\Delta\mathbf{s}_{i},\qquad\mathbf{q}_{i}^{(k)}=\Delta\mathbf{q}_{i}\otimes\mathbf{q}_{i},$ (6)
where $\otimes$ denotes multiplication of quaternions. Normalizing $\Delta\boldsymbol{q}_{i}$ produces a unit quaternion, so $\boldsymbol{q}_{i}^{(k)}$ remains a valid orientation. Eq. (6) therefore converts the canonical Gaussians $\{\boldsymbol{\mu}_{i},\boldsymbol{s}_{i},\boldsymbol{q}_{i}\}$ into the state-specific set $\{\boldsymbol{\mu}_{i}^{(k)},\boldsymbol{s}_{i}^{(k)},\boldsymbol{q}_{i}^{(k)}\}$ that describes the interaction state $k$.
For the $K$ observed interaction states of the object, we jointly optimize the network parameters $\theta_{f_{\text{def}}}$, the latents $\{\boldsymbol{z}_{k}\}_{k=1}^{K}$, and the standard Gaussian parameters by minimizing the results.
$\mathcal{L}_{\text{total}}=\underbrace{\sum_{k,v}\!\sum_{\mathbf{x}\in\Omega_{k}^{(v)}}\lVert\mathbf{C}_{k}^{(v)}(\mathbf{x})-\mathbf{I}_{k}^{(v)}(\mathbf{x})\rVert_{1}}_{\mathcal{L}_{\text{photo}}}+\mathcal{L}_{\text{D\_SIMM}},$ (7)
where $\mathbf{C}_{k}^{(v)}$ is the color rendered in pixel $\mathbf{x}$, $\mathbf{I}_{k}^{(v)}$ is the corresponding ground truth image, and $\mathcal{L}_{\text{D\_SIMM}}$ is the standard 3DGS density similarity term that discourages overlapping Gaussians with dissimilar colors.
After convergence, every latent $\boldsymbol{\alpha}_{k}$ produces a state-specific geometry $\{\boldsymbol{\mu}_{i}^{(k)},\boldsymbol{s}_{i}^{(k)},\boldsymbol{q}_{i}^{(k)}\}_{i=1}^{N}$ while sharing the appearance attributes $\{\mathbf{c}_{i},\alpha_{i}\}_{i=1}^{N}$. A single model therefore yields a coherent Gaussian scene for each interaction state and jointly encodes geometry, appearance, and articulation.
3.3 Coarse Part-Level Segmentation of Gaussian Primitives
Motion-driven separation of static and dynamic Gaussians.
Without object masks or semantic priors, we detect dynamic primitives by computing the maximum Euclidean displacement of each Gaussian center across the $K$ interaction states. Let $\boldsymbol{\mu}_{i}^{(k)}\in\mathbb{R}^{3}$ be the center of Gaussian primitives $g_{i}$ in state $k$. We define its maximal displacement is
$d_{i}=\max_{j,k\in\{1,\dots,K\}}\bigl{\lVert}\boldsymbol{\mu}_{i}^{(j)}-\boldsymbol{\mu}_{i}^{(k)}\bigr{\rVert}_{2},\quad\hat{d}_{i}=d_{i}/\!\max_{r}d_{r}\in[0,1].$ (8)
Gaussians with $\hat{d}_{i}\geq\tau_{\text{mot}}$ form the dynamic set $\mathcal{G}_{\text{dyn}}$, and the remainder $\mathcal{G}_{\text{stat}}$ is static, where the threshold $\tau_{\text{mot}}\in[0,1]$ specifies the fraction of the maximum displacement of the entire scene that qualifies as motion.
Estimating the number of motion parts via a VLM.
To infer how many rigid parts appear in $\mathcal{G}_{\text{dyn}}$ we sample $M$ image pairs $(I_{k_{i}}^{(v_{p})},I_{k_{j}}^{(v_{q})})$ with different states $k_{i}\neq k_{j}$. A visual language model (VLM) [39] is queried with the fixed prompt:
“Compare the two images. How many components moved?”
“Answer: ‘Number of moved components: [N]’.”
The number of parts $n_{\text{parts}}$ is set in the mode of the VLM predictions $M$, i.e., the integer that occurs most frequently, making the estimate insensitive to occasional miscounts.
Trajectory-based clustering of dynamic Gaussians.
For each $g_{i}\!\in\!\mathcal{G}_{\text{dyn}}$ we assemble a motion descriptor
$\mathbf{f}_{i}=\bigl{[}\hat{\Delta}\boldsymbol{\mu}_{i}^{(1)},\,\lVert\Delta\boldsymbol{\mu}_{i}^{(1)}\rVert_{2},\,\dots,\,\hat{\Delta}\boldsymbol{\mu}_{i}^{(K-1)},\,\lVert\Delta\boldsymbol{\mu}_{i}^{(K-1)}\rVert_{2}\bigr{]}\in\mathbb{R}^{4(K-1)},$ (9)
where
$\Delta\boldsymbol{\mu}_{i}^{(k)}=\boldsymbol{\mu}_{i}^{(k+1)}-\boldsymbol{\mu}_{i}^{(k)},\qquad\hat{\Delta}\boldsymbol{\mu}_{i}^{(k)}=\frac{\Delta\boldsymbol{\mu}_{i}^{(k)}}{\lVert\Delta\boldsymbol{\mu}_{i}^{(k)}\rVert_{2}+\varepsilon}.$ (10)
Unit directions encode orientation and lengths keep relative travel, so the descriptor is translation/rotation aware yet scale balanced. After normalization of $\ell_{2}$, we cluster the descriptors with K-means [40] using $K=n_{\text{parts}}$. On the unit sphere, Euclidean distance equals cosine distance, therefore Gaussians that share one rigid motion, even at different amplitudes, are grouped together.
Clusters containing fewer than $2\%$ of dynamic Gaussians are regarded noise and merged into the nearest larger cluster. Each dynamic primitive then receives a part label $c_{i}\in\{1,\dots,n_{\text{parts}}\}$, while static ones are labeled 00. These coarse labels initialize the fine-grained stage that follows.
3.4 Refined Part-Level Segmentation of Gaussian Primitives
Visibility-aware prompt generation for SAM.
To sharpen the coarse part labels in Sec. 3.3 we first render each part through a visibility filter and then draw sparse point prompts for the Segment Anything Model (SAM) [32].
Per-Gaussian contribution at a pixel. For a pixel $\mathbf{x}$ in view $v$, the contribution weight $w_{i}^{(v)}(\mathbf{x})$ from Gaussian $g_{i}=(\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i},\mathbf{c}_{i},\alpha_{i})$ is obtained by first evaluating its screen-space density $\rho_{i}^{(v)}(\mathbf{x})$ (Eq. (2)) and then applying the front-to-back compositing rule (Eq. (3)), which together map every three-dimensional primitive to a continuous weight field whose value at $\mathbf{x}$ equals $w_{i}^{(v)}(\mathbf{x})$.
Part specific visibility confidence.  For each pixel we accumulate the weight of part $p$ as $w_{p}^{(v)}(\mathbf{x})=\sum_{g_{i}\in\mathcal{G}_{p}}w_{i}^{(v)}(\mathbf{x})$ and compare it with the strongest competitor $w_{\max,\urcorner{p}}^{(v)}(\mathbf{x})=\max_{q\neq p}w_{q}^{(v)}(\mathbf{x})$. The pixel is assigned to part $p$ when
${w_{p}^{(v)}(\mathbf{x})}/\{w_{\max,\urcorner{p}}^{(v)}(\mathbf{x})+\varepsilon\}>\tau_{\text{vis}},$ (11)
where $\tau_{\text{vis}}=1.2$ and $\varepsilon$ avoids division by zero. The winner margin test keeps only pixels in which part $p$ clearly dominates, thus reducing the ambiguity at overlaps.
Sampling prompts. We define the set of visible pixels corresponding to part $p$ in view $v$ as $\Omega^{(v)}_{p^{+}}=\{\mathbf{x}\mid\text{pixel assigned to part }p\}$ and the non-contributing set as $\Omega^{(v)}_{p^{-}}=\{\mathbf{x}\mid w_{p}^{(v)}(\mathbf{x})=0\}$. Furthest point sampling (FPS) on $\Omega^{(v)}_{p^{+}}$ and $\Omega^{(v)}_{p^{-}}$ selects ten positive prompts $\mathcal{P}^{(v)}_{p^{+}}$ and twenty negative prompts $\mathcal{P}^{(v)}_{p^{-}}$, which together with the RGB image are fed to SAM to obtain the mask $\mathcal{M}_{p}^{(v)}$.
Boundary-aware Gaussian splitting.
For each Gaussian $g_{i}$ we find its nearest view $v_{\text{near}}$ by ray sampling and project the ellipsoid to a two-dimensional ellipse. If the ellipse major axis extends beyond the mask of the part $\mathcal{M}_{p}^{(v_{\text{near}})}$, the Gaussian becomes a boundary candidate.
We follow SAGD [36] to split each candidate, as shown in the bottom part of Fig. 1. The fraction of the major axis that lies inside the mask is measured in image space and the same ratio is applied to the center $\boldsymbol{\mu}_{i}$ and the scale vector $\mathbf{s}_{i}$, resulting in two children $g_{i}^{(\text{part})}$ and $g_{i}^{(\text{bg})}$. The part child retains the in-mask portion and the background child captures the overflow. The background child $g_{i}^{(\text{bg})}$ is re-evaluated in its nearest view; if its ellipse still crosses another part mask the split is applied recursively until every descendant lies inside exactly one mask.
Part-aware Gaussian field via local refinement.
Finally, parameters of unsplit Gaussians remain fixed, whereas all new children are locally fine-tuned for a few iterations on $\mathcal{L}_{\text{total}}$ (updating $\mathbf{q}$ and $\alpha$ only) to restore the photo consistency. The recursive split-and-tune process produces sharply aligned, mask-consistent part boundaries while leaving the converged Gaussian field undisturbed.
The entire procedure yields a refined, part-aware Gaussian field with inteactive state
$\mathcal{G}^{(\boldsymbol{\alpha})}=\bigl{\{}\mathcal{G}_{p}^{(\boldsymbol{\alpha})}\bigr{\}}_{p=1}^{n_{\text{parts}}},\qquad\mathcal{G}_{p}^{(\boldsymbol{\alpha})}=\bigl{\{}(\boldsymbol{\mu}_{i}^{(\boldsymbol{\alpha})},\boldsymbol{\Sigma}_{i}^{(\boldsymbol{\alpha})},\mathbf{c}_{i},\alpha_{i})\bigr{\}}_{i\in\mathcal{G}_{p}},$ (12)
where $\boldsymbol{\alpha}\in\mathbb{A}$ indexes the interaction state ($|\mathbb{A}|=K$) and $\mathcal{G}_{p}$ denotes the set of Gaussian indices belonging to part $p$. Each state-specific field $\mathcal{G}^{(\boldsymbol{\alpha})}$ encodes a coherent multi-part configuration, while the union over all $\boldsymbol{\alpha}$ compactly describes the full articulated motion range of the object.
3.5 Multi-Task Modeling for Articulated Objects
Given the part-aware Gaussian field $\mathcal{G}^{(\boldsymbol{\alpha})}=\{\mathcal{G}_{p}^{(\boldsymbol{\alpha})}\}_{p=1}^{n_{\text{parts}}}$, we extract three part-level outputs. First, every subset $\mathcal{G}_{p}^{(\boldsymbol{\alpha})}$ is rendered in depth and color images in all calibrated views; marching cubes applied to the resulting volume produce the mesh $\mathcal{M}_{p}\in\mathbb{R}^{|V_{p}|\times 3}$. Second, to classify the type of joint part $p$, we align two interaction states with the Kabsch algorithm [41] and inspect the residual displacement: a low-rank residual indicates a revolute joint, while a full-rank residual indicates a prismatic joint. Third, the trajectories of the Gaussian centroids belonging to part $p$ are fitted with a minimal motion model. A revolute joint is parameterised by a pivot point $\mathbf{p}\in\mathbb{R}^{3}$ and a unit quaternion $\mathbf{q}\in\mathbb{R}^{4}$ with $\lVert\mathbf{q}\rVert=1$. A prismatic joint is parameterised by a unit slide axis $\mathbf{a}\in\mathbb{R}^{3}$ with $\lVert\mathbf{a}\rVert=1$ and a translation distance $d\in\mathbb{R}$. More details are provided in Sec. A of the Appendix.
4 The RS-Art Dataset
We establish RS-Art, a high-quality, multi-modal benchmark that bridges the gap between synthetic and real-scene evaluation for articulated-object modeling. It comprises real-world captures of articulated objects alongside their reverse-engineered, part-level digital counterparts, providing a rare combination of dense sensory data and precise structural ground truth. Covering six representative object categories (drawers, desk lamps, eyeglasses, floppy-disk drives, woven baskets, and phone/laptop stands), the dataset includes three diverse instances per category. Each object is recorded in seven distinct articulation states, spanning both compound multi-joint configurations and single-joint extrema, resulting in over 400 wide-baseline RGB-D observations per instance. For every object, we reconstruct a high-fidelity, textured, multi-part mesh annotated with joint axes and motion limits. All assets are released in URDF format, enabling direct deployment in physics-based simulators for downstream tasks and rigorous, physically grounded evaluation. Further details of the dataset are provided in Sec. B of the Appendix.
5 Experiments
Table 1: Results on the Partnet-Mobility Dataset. A superscript denotes that the method failed to obtain a valid result for at least one object in the corresponding category.
Metric Method Manual Synthetic Objects Mean
Box Door Eyeglasses Faucet Oven Fridge Storage Table
Axis $\downarrow$ Ang 0 PARIS [20] 36.27 24.83 64.35 13.75 54.88 69.12 27.05 15.60 38.23
ArticulatedGS [21] 0.98 1.43 2.25 5.14 4.13 14.14 23.24 1.35 6.58
DTArt [23] 20.65 0.38 18.18 1.06 3.37 40.99 0.46 0.40 10.69
ArtGS [24] 0.01 0.35 0.06 8.05 2.33 0.13 6.05 1.98 2.37
Ours 0.49 0.31 0.06 0.88 0.34 0.09 0.21 0.25 0.33
[…]
ArticulatedGS [21] 5.12 0.77 2.12 4.12 5.25 2.13 0.12 0.52 2.52
DTArt [23] 42.60 0.75 1.62 1.44 14.21 2.17 0.22 18.12 10.14
ArtGS [24] 1.36 0.68 0.08 17.02 9.24 1.41 0.01 7.66 4.68
Ours 0.28 0.49 1.22 0.58 0.12 0.73 0.04 0.28 0.47
[…]
ArticulatedGS [21] 0.97 0.26 2.11 0.24 3.46 0.26 0.74 0.79 1.10
DTArt [23] 7.09 0.01 4.96 0.02 14.89 11.35 0.07 0.03 4.80
ArtGS [24] 0.06 0.03 0.06 0.01 1.9 0.02 1.28 39.14 5.31
Ours 0.06 0.02 0.02 0.04 0.27 0.16 0.03 0.01 0.08
[…]
ArticulatedGS [21] 0.24 0.62 0.24 1.24 2.15 0.15 0.24 0.65 0.69
DTArt [23] 3.96 0.21 11.75 0.02 32.87 0.15 0.01 0.01 6.12
ArtGS [24] 0.51 0.03 0.04 0.05 6.24 0.04 0.01 5.98 1.61
Ours 0.00 0.02 0.09 0.04 0.08 0.03 0.00 0.01 0.03
[…]
ArticulatedGS [21] 4.24 2.65 34.24 2.46 6.35 3.35 2.33 0.36 7.00
DTArt [23] 42.85 0.35 13.91 0.07 70.02 36.73 0.07 0.55 20.57
ArtGS [24] 2.78 0.05 0.07 6.80 1.67 0.03 6.69 8.97 3.38
Ours 2.09 0.05 0.02 0.95 0.74 0.10 0.43 0.32 0.59
[…]
ArticulatedGS [21] 20.04 7.24 0.63 9.24 2.58 19.87 0.90 0.23 7.59
DTArt [23] 63.48 89.31 104.92 1.44 65.79 4.91 0.05 0.21 41.26
ArtGS [24] 15.68 0.05 0.08 16.89 3.87 0.03 0.16 24.09 7.60
Ours 0.87 0.05 0.02 0.50 0.34 0.02 0.21 0.06 0.26
CD-s $\downarrow$ PARIS [20] 6.93 17.90 23.35 9.36 13.45 18.96 10.44 3.45 12.98
ArticulatedGS [21] 18.17 167.53 15.14 0.97 29.43 23.80 16.18 5.92 34.64
DTArt [23] 2.45 20.60 49.28 0.49 2.60 4.06 1.27 8.65 11.17
ArtGS [24] 3.48 0.35 0.13 1.05 5.44 9.69 5.10 9.04 4.29
Ours 1.97 0.80 0.11 0.64 1.66 4.15 2.78 4.27 2.05
CD-m 0 $\downarrow$ PARIS [20] 82.05 42.22 86.37 34.79 107.25 212.08 182.85 16.58 95.52
ArticulatedGS [21] 9.65 0.84 26.88 35.63 1.91 11.24 7.32 7.24 12.59
DTArt [23] 363.75 0.22 47.24 0.22 157.26 0.63 0.23 46.30 76.98
ArtGS [24] 2.47 0.44 0.16 0.38 8.89 0.86 3.46 143.78 20.06
Ours 2.44 0.15 0.09 0.71 1.79 0.82 0.03 5.40 1.43
CD-m 1 $\downarrow$ PARIS [20] 128.58 32.08 98.54 48.86 301.18 76.82 125.06 76.18 110.91
ArticulatedGS [21] 8.78 0.44 13.47 0.27 10.92 1.15 179.27 12.82 28.39
DTArt [23] 198.70 1.17 1133.05 0.22 140.52 0.36 0.27 0.22 184.31
ArtGS [24] 3.35 0.42 0.17 0.46 8.47 0.86 2.25 1145.91 145.23
Ours 1.84 0.43 0.13 0.74 2.15 0.27 0.11 0.76 0.80
CD-w $\downarrow$ PARIS [20] 6.30 12.59 17.24 9.98 10.93 18.89 8.86 2.81 10.95
ArticulatedGS [21] 13.75 12.59 19.03 2.78 23.22 18.89 50.59 5.81 18.33
DTArt [23] 1.08 1.40 0.15 0.50 1.73 0.98 10.17 1.39 2.18
ArtGS [24] 5.91 1.56 0.17 0.58 7.35 16.82 8.66 12.73 6.72
Ours 4.07 1.24 0.13 0.47 1.48 10.16 5.48 1.14 3.02
5.1 Experimental Setting
Datasets.
Most studies evaluate only one object per category, which limits the statistical power. We extend PartNet-Mobility [1] by selecting eight categories and two to three instances per category, each with two movable parts, so all methods face the same kinematic complexity. We also include four objects with three movable parts to test the robustness under higher joint counts. Real-scene performance is measured on the RS-Art we proposed. Together, the enlarged PartNet-Mobility split and RS-Art form a balanced test bed for synthetic and real scenarios.
Baselines.
We compare our method with recent SOTA methods. The single-joint baselines, which can process only one movable part at a time, are PARIS [20] and ArticulatedGS [21]. As ArticulatedGS [21] has no public code, we reproduced it with the official 3DGS implementation and the details described in the article. Each single-joint pipeline was run once for every joint: in each run, the input images show motion of a single joint while the other parts remain fixed, and the per-joint outputs were merged for evaluation. The multi-joint baselines, which can handle several movable parts in one pass, are DTArt [23] and ArtGS [24]. Both baselines require a prespecified part count, whereas our method infers the parts automatically.
Metrics.
We adopt exactly the evaluation metrics defined in DTArt [23], covering the geometry, part segmentation and precision of the joint parameter. The only extension is that our reported scores are aggregated over all movable parts of an object, rather than being limited to a single target part.
Additional implementation details are in Sec. C of the Appendix.
Figure 2: Qualitative multi-task modeling results on multi-part articulated objects.
5.2 Comparison on Objects with Multiple Movable Parts
Tab. 1 reports quantitative results on our expanded PartNet-Mobility split, which contains objects with two independently actuated parts. Previous studies usually average across several runs, yet we observe large run-to-run variance in several baselines, to reveal this sensitivity, we list the first run for every method. These numbers are not cherry-picked, and subsequent runs often produce even lower scores. DeGSS needs no manual part specification and still outperforms the baselines in most metrics. Fig. 2, which also shows an object with three movable parts, qualitatively confirms that our joint estimates align more closely with the ground truth. DTArt [23] benefits from depth supervision and therefore renders smoother surfaces, but still lags behind in articulation accuracy. Overall, the use of a latent-conditioned deformation field spanning multiple interaction states yields markedly higher stability than single-state or pairwise pipelines. Results on more objects with three movable parts are presented in Sec. D.1 of the Appendix, where the same trend persists.
Figure 3: Generalization to unseen interaction states. We interpolate the latent code while deforming only the Gaussians that belong to a chosen part.
5.3 Generalization to Unseen Interaction States
Fig. 3 illustrates that our method is the only one that is capable of disentangling and controlling the motion of several parts independently while maintaining plausible geometry in interaction states never shown during training. By interpolating the latent code and deforming only the Gaussians that belong to a chosen part, we generate smooth, collision-free trajectories for every drawer and for the door. In contrast, PARIS [20] cannot keep the parts separated and often collapses or blends them together. The capacity to vary multiple joints independently and to synthesize genuinely novel poses represents a substantial step toward high-fidelity digital twin modeling of articulated objects.
5.4 Result on the RS-Art Dataset
Fig. 4 shows qualitative results on a real object from our RS-Art dataset. Despite sensor noise and challenging lighting, the method reconstructs detailed part-level geometry and texture and reliably extrapolates to interaction states not seen during training. These observations indicate that our method remains stable on real data. Further results are provided in Sec. D.2 of the Appendix.
Figure 4: Real object validation on RS-Art.
5.5 Ablation Studies
The most informative ablation is to remove the coarse-to-fine refinement stage introduced in Sec. 3.3 and Sec. 3.4. Fig. 5 contrasts the results without of the refinement stage (top) with those with of it (bottom) in a cabinet and fridge that contain multiple movable panels. Without refinement, the reconstructed parts bleed into one another and the contact regions exhibit severe artifacts; after refinement, the mesh shows clean separation, sharper boundaries, and no interpenetration. The improvement stems from visibility-guided SAM prompting and boundary-aware Gaussian splitting, which force the Gaussian field to conform to the per-part masks. We observe the same trend across other objects and provide additional visual and quantitative comparisons in Sec. D.3 of the Appendix.
Figure 5: Ablation on the coarse–to–fine refinement stage. For the variant without refinement we reconstruct multiple parts directly from the K-means clusters obtained in the coarse stage.
6 Conclusion
We introduce DeGSS, the fully self-supervised framework for articulated object modeling within a unified 3D Gaussian Splatting paradigm. A latent-conditioned deformation network warps a standard Gaussian field across interaction states, and a visibility-guided coarse-to-fine scheme refines deformation trajectories into part-aware Gaussian clusters. The resulting field supports part-level mesh extraction, joint typing, and motion parameter estimation. In an expanded PartNet-Mobility split and the novel RS-Art dataset, DeGSS surpasses existing methods without requiring part count in advance, and its learned deformation space generalizes smoothly to unseen interaction states, a property well suited to digital twin scenarios. The present implementation assumes accurate camera poses and cannot reconstruct regions that remain fully occluded in all views, leaving robustness to pose noise and reasoning about unseen structures as directions for future work.
Appendix
The supplementary material is organized as follows. Sec. A elaborates on the technical details; Sec. B details the acquisition and reverse modeling pipeline used to construct the RS-Art dataset; Sec. C lists the complete experimental settings; Sec. D presents additional quantitative and qualitative results; and Sec. E discusses the current limitations of our approach and directions for future work. All code and datasets will be released upon publication.
Appendix A Method Details
A.1 Boundary-Aware Gaussian Splitting
Selecting boundary candidates.
For each 3D Gaussian primitive $g_{i}=(\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i},\mathbf{c}_{i},\alpha_{i})$, we first locate its ray–sampling nearest view $v_{\text{near}}$. Assuming a local affine projection, the covariance $\boldsymbol{\Sigma}_{i}$ is mapped to the image plane by
$\boldsymbol{\Sigma}^{\prime}_{i}\;=\;\mathbf{J}\mathbf{W}\boldsymbol{\Sigma}_{i}\mathbf{W}^{\!\top}\mathbf{J}^{\!\top},$ (13)
where $\mathbf{W}$ and $\mathbf{J}$ denote the camera projection matrix and its Jacobian, respectively [19]. The resulting $2\times 2$ matrix $\boldsymbol{\Sigma}^{\prime}_{i}$ defines an ellipse; its major axis is aligned with the eigenvector of $\boldsymbol{\Sigma}^{\prime}_{i}$ that has the largest eigenvalue. If either end of this axis lies outside the part mask $\mathcal{M}^{(v_{\text{near}})}_{p}$, the primitive is marked as boundary candidate.
Geometric split ratio.
Let $A^{\ast}$ and $B^{\ast}$ be the two end points of the projected major axis and let $O^{\ast}$ be their intersection with the mask boundary ($A^{\ast}$ lies inside, $B^{\ast}$ outside). Following the split strategy of Hu et al. [36], the in-mask proportion in image space is
$\lambda_{2\text{D}}\;=\;\frac{\|O^{\ast}A^{\ast}\|}{\|A^{\ast}B^{\ast}\|}\;.$ (14)
Because the projection is affine locally, the same ratio applies in 3D. With $\mathbf{e}$ denoting the unit eigenvector of $\boldsymbol{\Sigma}_{i}$ that corresponds to its largest eigenvalue and $s_{i}$ the length of the associated axis, the two children are
  • Part child: $\boldsymbol{\mu}_{i}^{\text{part}}=\boldsymbol{\mu}_{i}+\tfrac{1-\lambda_{2\text{D}}}{2}\,s_{i}\mathbf{e},\;s_{i}^{\text{part}}=\lambda_{2\text{D}}\,s_{i};$
  • Background child: $\boldsymbol{\mu}_{i}^{\text{bg}}=\boldsymbol{\mu}_{i}-\tfrac{\lambda_{2\text{D}}}{2}\,s_{i}\mathbf{e},\;s_{i}^{\text{bg}}=(1-\lambda_{2\text{D}})\,s_{i}.$
Both children inherit color, opacity, and spherical harmonic coefficients to maintain visual consistency.
Recursive splitting and mask consistency.
The background child is re-evaluated in its own nearest view; if its projected ellipse still crosses another part mask, the split is applied recursively until every descendant ellipse is fully contained in a single mask. The recursion stops when the ellipse lies entirely inside its mask or the new scale falls below a minimum threshold. After each split, the children are queued in descending order of projected error, so the most inaccurate primitives are processed first. Finally, a multiview voting scheme assigns a consistent mask label to every new Gaussian, ensuring cross-view agreement.
A.2 Mesh Extraction from the Part-Aware Gaussian Field
Given the part-aware Gaussian field $\mathcal{G}^{(\boldsymbol{\alpha})}=\{\,\mathcal{G}^{(\boldsymbol{\alpha})}_{p}\,\}_{p=1}^{n_{\text{parts}}}$ for the interaction state $\boldsymbol{\alpha}$, we recover explicit geometry by multiview depth rendering followed by implicit–to–explicit surface reconstruction. For each subset of parts $\mathcal{G}^{(\boldsymbol{\alpha})}_{p}$, we render a depth map in every calibrated view, treating each Gaussian primitive as an ellipsoid whose pose and covariance are projected onto the image plane using the standard 3DGS rasterizer. The rendered depths are fused with a truncated signed-distance function (TSDF); voxel size and the truncation threshold are chosen empirically to balance detail and noise. The resulting TSDF volume provides an implicit surface for part $p$, from which we extract a triangle mesh $\mathbf{M}_{p}\!\in\!\mathbb{R}^{|V_{p}|\times 3}$ using the Marching Cubes algorithm implemented in Open3D [42]. This procedure yields a high-resolution, watertight mesh for every rigid part, faithfully capturing its geometry while preserving the spatial correspondence with the underlying Gaussian field.
A.3 Joint Type Identification
For a part $p$ we gather Gaussian centroids in two states $P=\{p_{i}\}$ and $Q=\{q_{i}\}$, and align them with the Kabsch algorithm [41], which yields rotation $\mathbf{R}$ and translation $\mathbf{t}$. The residuals $d_{i}=q_{i}-(\mathbf{R}p_{i}+\mathbf{t})$ are stacked in the $3\times n$ matrix $\mathbf{D}=[\,d_{1}\;\cdots\;d_{n}]$, where the $n$ is the number of Gaussian kernels.
A motion of revolute leaves points on a common axis, so $\mathbf{D}$ is rank-1; a motion of prismatic rigidly translates the entire part, giving full rank. With singular values $\sigma_{1}\geq\sigma_{2}\geq\sigma_{3}$, we compute $r=({\sigma_{2}+\sigma_{3}})/{\sigma_{1}}.$ We label the joint revolute if $r<0.05$ and prismatic otherwise, a rule that proved to be reliable across all reference objects.
A.4 Joint Parameter Estimation
Revolute joint.
A revolute joint is parameterized by a pivot point $\mathbf{p}\!\in\!\mathbb{R}^{3}$ and a unit quaternion $\mathbf{q}\!\in\!\mathbb{R}^{4}$ with $\lVert\mathbf{q}\rVert=1$. From the Kabsch alignment [41] we obtain a rotation matrix $\mathbf{R}\!\in\!\mathbb{R}^{3\times 3}$. Its rotation axis is the normalized eigenvector $\mathbf{u}$ that satisfies $\mathbf{R}\mathbf{u}=\mathbf{u}$, and the rotation angle is $\theta=\arccos\bigl{(}(\operatorname{tr}(\mathbf{R})-1)/2\bigr{)}$. The pivot $\mathbf{p}$ is found by least-squares fitting the axis to the static centroids $\mathcal{C}_{\text{static}}$,
$\mathbf{p}=\operatorname*{arg\,min}_{\mathbf{x}}\sum_{c\in\mathcal{C}_{\text{static}}}\!\lVert(\mathbf{x}-c)\times\mathbf{u}\rVert^{2}.$ (15)
Finally, the quaternion becomes $\mathbf{q}=\bigl{(}\cos\frac{\theta}{2},\,\mathbf{u}\sin\frac{\theta}{2}\bigr{)}$.
Prismatic joint.
A prismatic joint is defined by a unit slide axis $\mathbf{a}\!\in\!\mathbb{R}^{3}$ and a translation distance $d$. Let $P=\{p_{i}\}$ and $Q=\{q_{i}\}$ be the part centroids in the two states. Their means are $\bar{p}=\tfrac{1}{n}\sum_{i}p_{i}$ and $\bar{q}=\tfrac{1}{n}\sum_{i}q_{i}$, so the rigid translation is $\mathbf{t}=\bar{q}-\bar{p}$. Hence $\mathbf{a}=\mathbf{t}/\lVert\mathbf{t}\rVert$ and $d=\lVert\mathbf{t}\rVert$. These parameters constitute the minimal motion model used in our evaluation.
Appendix B RS-Art Dataset
Figure 6: Capture platform.
Capture platform.
[…]
The operator mounts an Intel RealSense RGB-D camera [43] rigidly on the lid of a laptop, then walks a circular trajectory whose radius is chosen in real time from the live video feed to keep the target object at the desired image scale. Each circuit produces a set of synchronized RGB and depth frames that share a consistent chroma-key background, which simplifies subsequent background removal and calibration.
Figure 7: Example of multi-state acquisition. Using a PartNet-Mobility object for illustration, we capture eight interaction states that span the full motion range of every joint. States 0–3 (top row) depict multi-part motion in which all movable components change pose simultaneously, whereas States 4–7 (bottom row) show the single-part mode where one component moves while the others remain fixed.
Real-scene acquisition pipeline.
As illustrated in Fig. 7, to keep the real data fully compatible with the experimental settings used for the PartNet-Mobility dataset [1], we record two complementary interaction groups for each object. (i) Multi-part sequences: all movable parts are actuated simultaneously, and the capture spans four evenly spaced configurations that bracket the mechanical limits, for example, fully closed, half-open, three-quarter open, and fully open. (ii) Single-part sequences: each rotational or translational part is moved while the others remain fixed; for every part, we acquire two extreme poses corresponding to its minimum and maximum joint limits.
For each interaction state, we sample RGB-D images along a circular path whose viewing elevation is uniformly chosen in the range $30^{\circ}\!\sim\!60^{\circ}$. The operator walks once around the object with an Intel RealSense camera [43] rigidly mounted on a laptop lid, keeping the target centered in the live preview.
Figure 8: Reverse modeling of a basket object.
Modeling and registration workflow.
All digital assets are created in Isaac Sim 4.2 [44], whose USD pipeline provides both photorealistic rendering and physically consistent simulation (see Fig. 8 for a representative modeling scene). We first reverse-model each real object with calipers, producing a watertight USD mesh; subparts are manually segmented according to functional affordances and joint boundaries. High-resolution textures and PBR materials are authored in Blender 4.2.0 [45] and imported into Isaac Sim without loss of visual fidelity. The physical properties - mass, inertia, and friction - are assigned following the official Omniverse workflow [46]. The joint pivots and motion limits are calibrated by system identification, ensuring that the articulation of the digital twin in Isaac Sim reproduces the joint behavior observed on the physical object. The resulting USD files provide a pixel-accurate appearance and dynamics-ready articulation, forming a reliable ground truth for both vision and physics experiments.
Figure 9: RS-Art Dataset directory layout.
Dataset structure.
Fig. 9 outlines the RS-Art hierarchy. The dataset contains 18 articulated objects drawn from six everyday categories:drawer, lamp, eyeglasses, floppy-disk drive, basket, and phone / laptop stand with three instances per class. For every object we supply the complete set of multi-view RGB-D captures for each recorded interaction state together with the shared camera intrinsics and the per-view extrinsics; a fully reverse-modeling digital twin comprising a USD file that bundles part meshes, textures, physics properties, and joint definitions, a matching lightweight URDF, and watertight PLY meshes for geometry evaluation; and a concise meta.json that enumerates all parts, joint types, axes, limits, and labels every state as either multi-part or single-part motion. This combination of real imagery and fully articulated digital twins enables training, quantitative evaluation, and sim-to-real studies for part-aware geometry reconstruction and joint parameter estimation.
Appendix C Experiments Settings
To provide the most comprehensive comparison to date, we evaluated all publicly available methods that match our task definition of multiview, multi-state articulated object reconstruction. Since competing pipelines differ in input assumptions and functional scope, we group them into three categories and tailor the data acquisition protocol accordingly, aiming for maximum fairness while respecting the design of each method.
Single-part assumptions (PARIS [20] and ArticulatedGS [21]).
These methods can reconstruct only one movable part at a time. For every movable part, we capture two interaction states in which that part moves while all other parts remain fixed, and we run the pipeline independently for each part. Each state is sampled with 100 random, calibrated RGB views, giving 200 images per part.
Multi-part, two-state methods (DTArt [23] and ArtGS[24]).
These pipelines accept two distinct object states and can handle several movable parts jointly, provided that the part count is known. We therefore collect two interaction states and sample 100 random views per state (total 200 RGB images). The ground-truth number of movable parts is passed to both methods. As DTArt requires depth input, depth maps for every view are rendered in Gazebo [47].
Our method (DeGSS).
DeGSS handles an arbitrary number of parts and interaction states without knowing the part count. To match the total image budget of competing methods, we sample four interaction states in which multiple parts move simultaneously and capture 50 random views per state, producing the same 200-image input. No depth or manual parts information is provided.
This protocol equalizes the per-run image budget to 200 views, making the inputs comparable across methods. Note, however, that single-part baselines must be executed once for the each movable part, so they ultimately consume $200\times n_{\text{parts}}$ images per object.
Table 2: Joint-level metrics for PartNet-Mobility objects that contain three movable parts. A dash “–” marks entries that are not applicable because the corresponding joint is prismatic and therefore has no pivot position.
Object Method Axis Ang 0\!$\downarrow$ Axis Ang 1\!$\downarrow$ Axis Ang 2\!$\downarrow$ Axis Pos 0\!$\downarrow$ Axis Pos 1\!$\downarrow$ Axis Pos 2\!$\downarrow$ Axis Mot 0\!$\downarrow$ Axis Mot 1\!$\downarrow$ Axis Mot 2\!$\downarrow$
Box 102373 PARIS [20] 9.96 29.61 0.47 0.13 0.31 0.30 30.23 111.21 33.72
ArticulatedGS [21] 0.52 6.53 3.62 0.83 8.32 0.42 5.23 2.13 14.43
DTArt [23] 88.32 62.54 64.68 1.99 5.56 6.79 99.97 67.17 82.00
ArtGS [24] 0.29 0.05 0.21 0.19 0.01 0.39 15.89 0.17 6.02
Ours 0.23 1.34 0.19 0.03 0.23 0.14 1.01 1.59 2.83
Storage 45194 PARIS [20] 64.47 59.08 38.69 0.29 0.13 116.73 109.94 0.32
ArticulatedGS [21] 20.43 4.12 2.32 1.31 0.21 13.11 3.12 1.31
DTArt [23] 27.60 0.20 78.01 4601.75 0.17 59.95 0.16 0.37
ArtGS [24] 54.17 1.68 0.75 0.29 0.08 96.54 1.66 0.02
Ours 9.82 1.56 0.56 0.27 0.03 9.79 1.45 0.07
Storage 45271 PARIS [20] 0.77 87.40 76.58 0.22 0.15 0.17 87.02
ArticulatedGS [21] 1.34 4.24 2.13 0.73 0.42 0.14 0.84
DTArt [23] 4.23 80.96 74.26 12.52 0.01 0.19 50.65
ArtGS [24] 0.02 0.85 0.03 0.14 0.03 0.01 0.02
Ours 0.00 0.34 0.26 0.03 0.00 0.00 0.58
Table 23372 PARIS [20] 1.69 89.40 79.91 0.52 0.84 0.18 157.14 109.55
ArticulatedGS [21] 0.62 5.31 6.24 5.31 0.52 0.54 4.24 13.12
DTArt [23] 0.38 0.22 0.19 0.21 0.01 0.04 0.14 0.48
ArtGS [24] 0.02 89.03 14.64 0.57 1.79 0.02 75.13 13.43
Ours 0.01 0.12 0.29 0.03 0.12 0.00 0.12 0.45
Table 3: Geometry-level metrics for PartNet-Mobility objects that contain three movable parts.
[…]
Box 102373 PARIS [20] 17.70 22.36 279.47 184.97 16.66
ArticulatedGS [21] 16.86 13.59 23.99 5.18 13.23
DTArt [23] 3.88 121.66 199.49 2.25 0.72
ArtGS [24] 9.70 14.18 5.38 2.63 7.78
Ours 6.70 0.88 0.39 2.21 5.75
Storage 45194 PARIS [20] 28.16 8.77 193.38 656.10 176.48
ArticulatedGS [21] 14.48 1.29 0.85 16.15 10.71
DTArt [23] 4.41 71.85 0.09 33.20 1.12
ArtGS [24] 50.88 19.00 1.78 2041.00 6.77
Ours 8.88 0.49 0.51 4.53 6.04
Storage 45271 PARIS [20] 5.09 668.39 703.55 19.35 5.50
ArticulatedGS [21] 14.48 1.28 0.85 16.15 10.71
DTArt [23] 2.94 307.65 F 406.23 8.76
ArtGS [24] 1.40 0.38 0.46 0.14 1.41
Ours 4.24 0.23 0.41 0.23 3.26
Table 23372 PARIS [20] 4.14 5.40 25.67 45.57 2.03
ArticulatedGS [21] 5.72 25.34 0.72 15.96 5.10
DTArt [23] 1.31 0.37 0.18 1.21 0.95
ArtGS [24] 3.38 4.50 1.34 1.05 2.84
Ours 2.17 1.19 0.15 0.25 4.19
Appendix D Additional Results
D.1 Results on Objects with Three Movable Parts
Sec. 5.2 in the main paper reported our main evaluation on objects that contain two movable parts. Here, we extend the analysis to the more challenging subset, objects equipped with three independently actuated components, and provide both quantitative and qualitative evidence of performance.
Tab. 2 and Tab. 3 compare all baselines in the three-part instances of PartNet-Mobility [1]. Our method achieves the lowest or second–lowest error in nearly every joint metric (axis angle, axis position, axis motion) and remains highly competitive in the geometry metrics (Chamfer distances). The consistently smaller joint errors indicate that the latent–conditioned deformation field captures the complex motion couplings that arise when three parts move independently, while the geometry scores confirm that the boundary regions between parts are reconstructed sharply.
Fig. 10 compares our method with the baselines on the objects with three movable parts and visualizes all recovered outputs: part-aware geometry (colored meshes), joint axes (red arrows) and part labels. The results demonstrate that our deformable-field formulation yields a coherent solution across all three modalities: the reconstructed geometry is complete and well aligned, the estimated axes coincide with the true joint locations and orientations, and the segmentation cleanly separates the three movable parts. Together, these visual cues indicate that the latent-conditioned deformation field is able to discover and model complex kinematic structures, even with three interacting joints, without manual specification of part identities.
Figure 10: Qualitative multi-task modeling results on articulated objects with three movable parts.
Table 4: Results on the RS-Art Dataset.
Metric Method Real-World Objects Mean
Drawers Desk Lamps Eyeglasses Floppy-Disk Drives Woven Baskets Phone/Laptop Stand
Axis $\downarrow$ Ang 0 ArtGS [24] 41.86 72.98 54.24 29.24 45.56 63.54 51.24
Ours 0.84 0.32 4.45 1.71 0.88 0.54 1.46
[…]
Ours 0.25 0.34 0.90 1.08 0.58 15.27 3.07
[…]
Ours 0.36 2.07 0.89 1.17 1.12
[…]
Ours 0.32 0.36 0.40 9.31 2.60
[…]
Ours 0.58 0.87 0.63 1.62 9.46 1.79 2.49
[…]
Ours 0.36 0.58 0.74 2.27 4.97 5.04 2.33
CD-s $\downarrow$ ArtGS [24] 41.39 11.38 25.24 53.53 6.34 14.35 25.37
Ours 13.13 6.25 10.19 26.88 4.05 13.36 12.31
CD-m 0 $\downarrow$ ArtGS [24] 121.69 39.09 33.24 104.35 58.24 26.49 63.85
Ours 3.09 4.83 7.19 10.37 5.63 10.08 6.86
CD-m 1 $\downarrow$ ArtGS [24] 180.57 34.18 234.24 156.35 62.64 43.98 118.66
Ours 8.57 3.86 9.41 13.47 0.27 41.44 12.84
CD-w $\downarrow$ ArtGS [24] 7.96 6.09 20.35 96.35 36.25 35.96 33.83
Ours 16.55 3.05 14.80 19.03 12.78 29.90 14.35
D.2 Result on the RS-Art Dataset
Fig. 11 and Tab. 4 present qualitative and quantitative results on our RS-Art dataset. Although the capture setup avoids strong specular highlights and cast shadows, the real objects still exhibit complex materials, weak surface texture, and a narrower viewing envelope than in simulation, making the task substantially harder than on PartNet-Mobility [1]. Among the baselines, only ArtGS can be made to run reliably on the real data, so we compare against it on every instance that yields a converged model. Across all evaluated objects, our method achieves lower joint parameter and geometry errors and produces cleaner part segmentation, confirming that the deformation-field formulation transfers to real scenes. Meanwhile, the results highlight the open challenges discussed in Sec. E: fine surface details degrade under limited texture and grazing angles, and heavy occlusion can still disturb the 3DGS reconstruction. These observations underline the necessity of RS-Art as a bridge between synthetic evaluation and real-world deployment.
Figure 11: Qualitative comparison on real objects from RS-Art.
Table 5: Quantitative ablation study on the refinement stage.
Object Method CD-$\downarrow$ CD-m 0$\downarrow$ CD-m 1$\downarrow$ CD-$\downarrow$
Box 100676 w/o Refined 11.66 2.12 2.14 10.16
w/ Refined 9.42 1.58 1.52 8.03
Oven 7187 w/o Refined 21.56 5.44 2.16 17.58
w/ Refined 16.58 0.32 0.54 13.91
Refrigerator 12248 w/o Refined 16.94 0.60 0.92 12.06
w/ Refined 13.63 0.21 0.23 9.53
Storage 47254 w/o Refined 13.63 2.91 1.24 9.53
w/ Refined 6.52 0.60 0.92 5.05
D.3 Ablation Studies
Quantitative impact of the refinement stage.
Sec. 5.5 in the main paper visualizes the qualitative benefit of coarse to fine refinement. Tab. 5 complements that visual evidence with numerical results on four representative objects. Across all geometry metrics, the refined model (w/ Refined) consistently outperforms the same network trained without the refinement stage (w/o Refined). Averaged across the four objects, the refinement reduces CD-s by $35\%$ and CD-m by $60\%$. The greatest gains occur at part boundaries (CD-m columns), confirming that boundary-aware Gaussian splitting and local fine-tuning markedly sharpen the reconstructed surfaces and prevent inter-penetration. These numbers corroborate the qualitative observations in Sec. 5.5 and highlight the refinement stage as a key component for accurate part-level geometry.
Table 6: Ablation on the number of interaction states $K$. Each column lists the performance metrics obtained with the corresponding $K$, evaluated on the PartNet-Mobility object Door 9168.
Metric $K{=}2$ $K{=}4$ $K{=}6$ $K{=}8$ $K{=}10$
Axis Ang 0$\downarrow$ 31.45 0.30 0.14 0.07 0.36
Axis Ang 1$\downarrow$ 47.83 0.33 0.22 0.17 0.15
Axis Pos 0$\downarrow$ 0.38 0.34 0.12 0.06 0.18
Axis Pos 1$\downarrow$ 0.07 0.26 0.19 0.05 0.29
Part Motion 0$\downarrow$ 57.60 1.44 1.35 1.24 1.03
Part Motion 1$\downarrow$ 56.71 2.38 1.74 0.71 2.10
CD-$\downarrow$ 1.63 0.53 0.56 0.63 0.77
CD-m 0$\downarrow$ 85.87 0.63 0.49 1.38 1.24
CD-m 1$\downarrow$ 142.74 0.33 0.57 0.90 0.79
CD-$\downarrow$ 0.81 0.60 0.64 0.84 0.87
Figure 12: Qualitative effect of the number of interaction states.
Impact of the number of interaction states.
Tab. 6 and Fig. 12 show how our method reacts to different numbers of input interaction states $K$. With just two interaction states ($K=2$), the network sees only a start-to-end displacement for each Gaussian. This limited evidence allows multiple trajectory assignments to fit the same images, so Gaussians can drift between neighboring parts, leading to inflated joint errors and indistinct part boundaries. As more states are provided, every primitive must follow a longer and more distinctive trajectory. The stronger constraint eliminates ambiguity: a Gaussian that does not truly belong to a part can no longer match the motion of that part in all states, and the optimizer moves it to the correct rigid component, causing a rapid drop in error. However, after each joint has appeared in several diverse poses, additional states add little fresh kinematic information while still introducing measurement noise and extra training time, so the accuracy curve levels off and may fluctuate slightly.
Appendix E Limitations
Incomplete part reconstruction under occlusion.
Like other NeRF- and 3D Gaussian Splatting-based approaches, DeGSS may struggle to reconstruct complete geometry and texture for every part of the object. This limitation arises in cases of severe occlusion, limited viewpoints, or coupled articulation, where certain surfaces are never observed. In addition, the representation is not surface-aware by design, with both the NeRF and 3DGS models radiance and density fields rather than explicit meshes—which can lead to surface artifacts and coarse or noisy geometry. Real-scene results also suggest that texture fidelity can degrade under challenging lighting conditions. Future work may explore integrating surface priors or reflectance models to improve reconstruction quality in under-constrained regions.
Limited articulation understanding from sparse states.
DeGSS infers motion by comparing interaction states, but reasoning about joint types and motion limits remains difficult. In particular, screw-type joints are difficult to identify, as the motion between observed states may produce little or no visible appearance change, making them indistinguishable from static or simpler joints under visual comparison alone. Similarly, estimating articulation limits (e.g., joint range or end stops) requires observing motions near those extremes, which sparse or biased sampling may not capture. Incorporating additional interaction sequences or learning from temporal transitions could help enrich articulation inference in future extensions.
Lack of physical properties for simulation.
While DeGSS focuses on visual and kinematic modeling, it does not capture physical properties such as mass, friction, or compliance, which are essential for high-fidelity digital twins. These properties are crucial for accurate simulation and downstream robotic tasks, but are currently absent from most NeRF/3DGS-based pipelines. Bridging this gap may require integrating material prediction models, learning from physical interactions, or coupling neural reconstruction with differentiable simulation frameworks.